# Challenges of Zero-Shot Recognition with Vision-Language Models:Granularity and Correctness

### v1

@@ Abstract
This paper investigates the challenges of applying vision-language models (VLMs) to zero-shot visual recognition tasks in an open-world setting, with a focus on contrastive vision-language models such as CLIP. We first examine the performance of VLMs on concepts of different granularity levels. We propose a way to fairly evaluate the performance discrepancy under two experimental setups and find that VLMs are better at recognizing fine-grained concepts. Furthermore, we find that the similarity scores from VLMs do not strictly reflect the correctness of the textual inputs given visual input. We propose an evaluation protocol to test our hypothesis that the scores can be biased towards more informative descriptions, and the nature of the similarity score between embedding makes it challenging for VLMs to recognize the correctness between similar but wrong descriptions. Our study highlights the challenges of using VLMs in open-world settings and suggests directions for future research to improve their zero-shot capabilities.
@@ 1 Introduction
Vision-language models (VLMs) have shown impressive capabilities in a wide range of tasks, including image captioning (Wang et al., 2022; Yu et al., 2022), visual question answering (Chen et al., 2023), and zero-shot visual recognition (Radford et al., 2021; Zhou et al., 2022; Gu et al., 2021; Xu et al., 2022). In particular, zero-shot visual recognition, the ability to recognize new, unseen objects or concepts without any additional training data, is an important area of research in the field of computer vision. Large-scale pre-trained VLMs, like CLIP (Radford et al., 2021), map visual inputs and language inputs to a shared latent space and output similarity/matching score between the aligned visual and language embeddings. Zero-shot visual recognition therefore can be achieved based on the the cross-modality score between visual inputs and the language defining any novel concept. For example, for zero-shot classification, the textual inputs can be prompts with class names e.g. “an image of flowers". Many VLMs trained on large scale image and text datasets (Jia et al., 2021; Yuan et al., 2021; Schuhmann et al., 2021, 2022a) reported promising zero-shot classification performance on existing benchmarks like ImageNet (Russakovsky et al., 2015) and multiple domain specific classification datasets. However, these benchmarks can hardly simulate the zero-shot recognition in a more realistic open-world setting. Therefore, the effectiveness of VLMs models on zero-shot recognition in open-world settings remains an open question.
In this paper, we focus on analyzing the characteristics of vision-language models, specifically contrastive models like CLIP, for open-world zero-shot capabilities. An ideal open-world zero-shot model, should recognize any language-defined input, e.g. simple words like the category names or complex descriptions like a caption, and output scores indicating whether the visual input correctly imply the semantics of the language input. Hence, we test two desired properties of open world zero-shot models: granularity and correctness.
First, VLMs should have consistent understanding of concepts at different granularity. Take an example shown in Figure 1-Left, if a model can successfully recognize an image of a leopard given the language input of a fine-grained concept, such as “an image of a leopard”, it should also recognize it with other valid descriptions of coarse-grained concepts, like “an image of a feline”. To fairly evaluate the consistency between concepts at different granularity, we measure the performance discrepancy of zero-shot classification for a coarse-grained class between predicting using the coarse-grained class name and aggregating the predictions using the names of its fine-grained children classes. Our experiments on ImageNet with two settings (two-level single-label classification and multi-label classification) both show that VLMs are rather better at recognizing fine-grained concepts than high-level coarse-grained concepts and the training data may account for the observed behavior. As we observe that training data sources of VLMs show significant impact on the measured performance discrepancy, we further analyze the distribution of concepts at different granularities in LAION dataset and find the moderatly more fine-grained concepts are more presented in image alt-text. The strong correlation between the frequency gap and performance discrepancy of ancestor labels and leaf children demonstrate the impact of data distribution.
Second, we study whether the cross-modality similarity scores from VLMs imply the correctness of language input to the visual input rather than only holistically similarity due to the contrastive learning objectives. We design hard positive text, e.g. single-label prompts with less information, that may have lower scores although being correct, as well as hard negative text, e.g. captions modified with errors, that may have higher scores although incorrect given the image. In the MS-COCO Lin et al. (2014) image-to-text retrieval task as shown in Figure 1-Right, retrieving hard positive text from hard negative text is extremely challenging, which suggests that the scores of VLMs cannot indicate correctness and are limited for open-world zero-shot recognition.
Our findings show that there is still a long way to use vision-language models for zero-shot visual recognition in the open-world setting. To our best knowledge, this is the first detailed study that evaluating VLMs for zero-shot recognition, in particular from the perspective of granularity and correctness. We hope our results can shed lights on better understanding the limitations of current VLMs. We also discuss potential directions for improving the issues from the angle of the training data and the model design, and connect to recent related research.
@@
Figure 1: Left: Zero-shot models should recognize images with fine-grained (FG) concepts such as “Leopard”, as well as coarse-grained (CG) concepts like “Feline” However, they often exhibit performance discrepancies on concepts at different levels of granularity. Right: Zero-shot models should recognize whether the text correctly describe the given image. However, constrastive vision-language modelsstruggle to recognize the correctness of challenging positive text like single-label prompts and hard negatives like poisoned captions with small changes.
@@ 2 Related Works
CLIP-like vision-language foundation models have enabled open vocabulary visual recognition by mapping images with their corresponding language descriptions. Early methods (Radford et al., 2021; Jia et al., 2021) demonstrate the effectiveness of this paradigm on the image classification tasks. For example, CLIP is able to achieve decent zero-shot accuracy on 27 image classification datasets. Given its potential, the language-driven visual recognition paradigm has been extended to tasks including object detection (Zhou et al., 2022), semantic segmentation (Xu et al., 2022), video action recognition (Wang et al., 2021), depth estimation (Zhang et al., 2022), etc. Such language guided visual recognition has become the new paradigm in the field of computer vision since it can recognition new objects without any training data. In this paper, we would like to stress test these VLMs in terms of zero-shot visual recognition to better understand their capability and limitation in realistic open-world setting.
Studies of vision-language models
Thanks to the larger datasets and larger transformer models, many powerful vision-language models have been developed and shown great capability (Yu et al., 2022; Wang et al., 2022; Chen et al., 2023). At the same time, these models are being studied from various perspectives, such as robustness, bias, and other limitations (Galindo & Faria, 2021; Fort, 2021; Goh et al., 2021; Noever & Noever, 2021; Daras & Dimakis, 2022).  Qiu et al. (2022) investigates the robustness of nine open-sourced image-text models under common perturbations on five tasks, while  Schiappa et al. (2022) studies the robustness of video-text models.  Fang et al. (2022) further analyzes the robustness of VLMs under challenging natural distribution shifts and show that the more diverse training distribution is the main cause for the robustness gains.  Yuksekgonul et al. (2022); Thrush et al. (2022) systematically evaluates the ability to encode compositional information of the VLMs.  Cho et al. (2022) investigates the visual reasoning capabilities and social biases of different text-to-image models. To improve transferability,  Shen et al. (2022) designs an efficient and scalable approach that leverages external knowledge to learn image representations. In this paper, we study VLMs from two new perspectives: granularity and correctness through the lens of zero-shot visual recognition.
3 Zero-shot Visual Recognition With Vision and Language Models
@@
Here, the $\odot$ operator computes the score between visual and language embeddings, with cosine similarity being the common choice. For classification tasks, $x_{t}$ can be a class prompt, such as ’a photo of a car," or it can incorporate additional class-specific knowledge to improve performance. In our subsequent studies, we adopt the prompt templates used in  Radford et al. (2021) for classification tasks. We simplify $E_{t}(x_{t})$ and $f(x_{v},x_{t})$ for a class $y$ to $E_{t}(y)$ and $f_{cls}(x_{v},y)$, respectively.

### v3

# Benchmarking Zero-Shot Recognition with Vision-Language Models: Challenges on Granularity and Specificity

@@ Abstract
This paper presents novel benchmarks for evaluating vision-language models (VLMs) in zero-shot recognition, focusing on granularity and specificity. Although VLMs excel in tasks like image captioning, they face challenges in open-world settings. Our benchmarks test VLMs’ consistency in understanding concepts across semantic granularity levels and their response to varying text specificity. Findings show that VLMs favor moderately fine-grained concepts and struggle with specificity, often misjudging texts that differ from their training data. Extensive evaluations reveal limitations in current VLMs, particularly in distinguishing between correct and subtly incorrect descriptions. While fine-tuning offers some improvements, it doesn’t fully address these issues, highlighting the need for VLMs with enhanced generalization capabilities for real-world applications. This study provides insights into VLM limitations and suggests directions for developing more robust models.
@@ 1 Introduction
Vision-language models (VLMs) have shown impressive capabilities in a wide range of tasks, including image captioning [29, 32], visual question answering [1], and notably, zero-shot visual recognition [18, 36, 8, 30]. Models pretrained on large-scale image-caption datasets  [10, 33, 22, 23], like CLIP [18], have been at the forefront. These models achieve this by mapping visual and linguistic inputs to a shared latent space, enabling the recognition of novel objects or concepts in a zero-shot manner—a capability critical for developing versatile and intelligent visual systems.
While current VLMs perform well in various tasks, their application in open-world scenarios poses unique challenges. An ideal open-world zero-shot model would recognize any language-defined input, from simple concepts like “an image of flowers" to more complex descriptions like “a person playing with a dog on the beach", and output scores indicating whether the visual input correctly implies the semantics of the language input. Existing works often evaluate the zero-shot capability on various classification dataset like ImageNet [20] and domain specific datasets [11] without the notion of granularity of concepts, as well as image and text retrieval on Flickr30K [15] and COCO [13] that are not able to reveal the general failure pattern. These benchmark fall short of replicating the complexities of a realistic open-world setting, leaving a substantial gap in our understanding of the effectiveness of VLMs in such scenarios.
@@
Figure 1: Left: Zero-shot models should recognize images with fine-grained (FG) concepts such as “Leopard”, as well as coarse-grained (CG) concepts like “Feline” However, they often exhibit performance discrepancies on concepts at different levels of granularity. Right: Zero-shot models should recognize whether the text correctly describe the given image. However, vision-language models could be sensitive to the specificity of text and struggle to distinguish between the challenging positive like single-label prompts and hard negatives like poisoned captions with small changes.
This paper present new benchmarks on the pivotal properties when deploying VLMs for real-world zero-shot recognition: granularity and specificity. The first benchmark examines VLMs’ ability to consistently understand concepts at different levels of granularity. For instance, a model should recognize an image of a leopard both when presented with a fine-grained query like “an image of a leopard” and a more coarse-grained query such as “an image of a feline.” This consistency is crucial, not just as an indicator of the model’s comprehension of concept relationships but also for practical applications. A pertinent example is in autonomous driving, where recognizing “road cone” but failing to identify “barrier” could be problematic. To assess this, we employ an evaluation protocol where we measure the performance discrepancy in recognizing a coarse-grained class, both by directly using the coarse-grained class prompt and by aggregating predictions from its fine-grained children classes. We leverage a dataset with hierarchical labels, adapting ImageNet and its semantic hierarchy from WordNet.
The second benchmark evaluates how the specificity of language inputs affects VLM outputs, even when visual and linguistic inputs align. For example, a simple prompt like “a picture with a dog” may receive a lower score compared to a more detailed but incorrect caption “a dog and cow lying together on an orange couch.” This distinction is key in determining whether VLMs can accurately reflect the correctness of the alignment between visual and language inputs, rather than just overall similarity. To test this, we use an image-to-text retrieval task on the MS-COCO dataset, designing challenging positive texts with varying levels of specificity, such as single-label prompts with limited information, and hard negative texts like slightly modified but incorrect captions.
Our carefully designed benchmarks led to an extensive evaluation of state-of-the-art vision-language models (VLMs). We focus on contrastive models like CLIP, covering various aspects, including pretraining datasets, architectural designs, cross-modality interactions, and learning objectives. We discovered that VLMs face significant challenges in both benchmarks.
In the granularity evaluation, we find that VLMs prefer moderately fine-grained concepts over more abstract, coarse-grained ones. This tendency seems closely tied to the nature of the training data. A detailed analysis of the LAION dataset revealed a higher presence of moderately fine-grained concepts in image alt-text, suggesting that data distribution plays a critical role. In the specificity evaluation, VLMs showed sensitivity to text specificity: texts that differ in specificity from the training data, such as straightforward single-label prompts or overly detailed captions, often received lower scores than more precisely detailed but slightly erroneous captions. This challenges the VLMs’ ability to accurately distinguish between correct and subtly incorrect descriptions, complicating the retrieval of hard positive texts from hard negatives. The implication is that VLMs’ scoring does not always reliably reflect the true alignment between visual and textual inputs. Our exploration into fine-tuning VLMs with these hard text samples revealed that while it offers some improvements, it does not fully resolve the challenges as a complete solution.
To our best knowledge, this is the first comprehensive study that evaluating VLMs from the perspective of semantic granularity and specificity. We believe that the carefully designed benchmark provide a valuable tool to the community to better quantitatively evaluate VLMs. With the proposed benchmark, we observed that all models surprisingly performs significantly worse than what we may hope. The findings and insights from our analysis may shed lights on better understanding the limitations of current VLMs and the challenges of using it for zero-shot recognition, and inspire new models with better generalization.
@@ 2 Related Works
CLIP-like vision-language foundation models have enabled open vocabulary visual recognition by mapping images with their corresponding language descriptions. Early methods [18, 10] demonstrate the effectiveness of this paradigm on the image classification tasks. For example, CLIP is able to achieve decent zero-shot accuracy on 27 image classification datasets. Given its potential, the language-driven visual recognition paradigm has been extended to tasks including object detection [36], semantic segmentation [30], video action recognition [28], depth estimation [35], etc. Such language guided visual recognition has become the new paradigm in the field of computer vision since it can recognition new objects without any training data. In this paper, we would like to stress test these VLMs in terms of zero-shot visual recognition to better understand their capability and limitation in realistic open-world setting.
Benchmarking vision-language models
Thanks to the larger datasets and larger transformer models, many powerful vision-language models have been developed and shown great capability [32, 29, 1]. At the same time, these models are being studied from various perspectives, such as robustness, bias, and other limitations [6, 5, 7, 14, 3].  [17] investigates the robustness of nine open-sourced image-text models under common perturbations on five tasks, while  [21] studies the robustness of video-text models.  [4] further analyzes the robustness of VLMs under challenging natural distribution shifts and show that the more diverse training distribution is the main cause for the robustness gains.  [34, 27] systematically evaluates the ability to encode compositional information of the VLMs.  [2] investigates the visual reasoning capabilities and social biases of different text-to-image models. To improve transferability,  [25] designs an efficient and scalable approach that leverages external knowledge to learn image representations. In this paper, we study VLMs from two new perspectives: granularity and specificity through the lens of zero-shot visual recognition.
Table 1: An overview of the differences between the vision-language models evaluated in our study by the architecture, pretraining datasets, learning objectives, and if using cross-modality fusion modules. ITC, ITM, MIM, MTM, MMM stands for image-text contrastive, image-text matchng, masked image modeling, masked text modeling and masked multimodal modeling losses.
Model Architecture Datasets Objectives Fusion
CLIP ViT-B-32 Private400M ITC -
ViT-L-14
OpenCLIP ViT-B-32 LAION400M ITC -
ViT-B-32 LAION2B
ViT-L-14
VIT-H-14
UniCL Swin-B YFCC14M ITC -
IN21K
IN21K+YFCC14M
IN21K+YFCC14M+GCC15M
KLITE Swin-B IN21K+YFCC14M+GCC15M ITC -
BLIP ViT-B-16 COCO+VG+CC+SBU +LAION+CapFilt-L ITC + ITM + Captioning -
$\checkmark$
FLAVA ViT-B/16 PMD70M
ITC+ITM
+MMM+MIM
+MTM
3 Zero-shot Visual Recognition With Vision-Language Models
@@
Here, the $\odot$ operator computes the score between visual and language embeddings, with cosine similarity being the common choice while some models like FLAVA use an additional module to fuse the multi-modal embeddings. For classification tasks, $x_{t}$ can be a class prompt, such as “a photo of a car", or it can incorporate additional class-specific knowledge to improve performance. In our subsequent studies, we adopt the prompt templates used in  [18] for classification tasks. We simplify $E_{t}(x_{t})$ and $f(x_{v},x_{t})$ for a class $y$ to $E_{t}(y)$ and $f_{cls}(x_{v},y)$, respectively.